OpenStudy (anonymous):

Please Help!!!! Find all solutions in the interval [0, 2π).? 1. 7 tan^3(x) - 21 tan x = 0 2. 2 sin^2(x) = sin x 3. Find all solutions to the equation. Please show your work. cos2x + 2 cos x + 1 = 0

OpenStudy (anonymous):

@rsnoob can u help?

OpenStudy (anonymous):

@laura*

OpenStudy (anonymous):

@sourwing can you please help me figure this out

OpenStudy (anonymous):

Problem 1 use algebra to rearrange and make into a simpler function. Then graph the function and find the x-intercepts. \[7\tan ^{3}x-21tanx=0\] \[7\tan ^{3}x=21tanx\] \[\tan ^{3}x=3tanx\] \[\frac{ \tan ^{3}x }{ 3tanx }=0\] \[\frac{ \tan ^{2}x }{ 3 }=0\] \[\tan ^{2}x=0\]

OpenStudy (anonymous):

You can do similar elimination of terms with the remaining equations.. For question 1 remember that the only time that tan^2 can be zero is then tanx is zero. Therefore your answer will be all places where tanx=0 on your interval (This should be a memorized graph for most Algebra II/Pre-Calc/Calc I classes).

OpenStudy (anonymous):

but these are my choices: pie/3 , 2pi/3, 4pi/3, 5pi/3 0, pi/5, π, 6pi/5 0, pi/3, 2pi/3, π, 4pi/3, 5pi/3 0, pi/3, π, 4pi/3

OpenStudy (anonymous):

The Original Function Graph is attached. So evaluating by looking at the graph would be the 3rd response. (I used google to get the graph)

OpenStudy (anonymous):

thankyou thankyou!! can you help with the others???

OpenStudy (anonymous):

I thought it was that but i wasn't quite sure

OpenStudy (anonymous):

2 * sin^2(x) = sin(x) dividing both sides of equation by sin(x) gets: (2* sin^2(x))/sin(x) = 1 dividing both sides of eqution by 2 gets: sin^2(x)/sin(x) = 1/2 since sin^2(x) = sin(x) * sin(x), equation becomes: (sin(x)*sin(x))/sin(x) = 1/2 which becomes: sin(x) = 1/2 ----- if sin(x) = 1/2, then x = 30 degrees. ----- since sin(x) = hypotenuse / y value on the graph, and since hypotenuse is always positive, and since y value on the graph is positive in quadrants 1 and 2, then sin (180-x) = sin(x). ----- since 180-30 = 150, then x can be either 30 degrees or 150 degrees. Ooops I forgot you were doing radians so just convert

OpenStudy (anonymous):

I will help with the third if you rewrite using the double angle identity for cos and try some of the elimination steps I have shown.

OpenStudy (anonymous):

I'm at cos^2x=-2cos-1